OpenStudy (anonymous):

Suppose you have a cookie stand, and when you charge $3 per cookie box you sell 200 boxes. But when you raise your price to $4 you only sell 120 boxes. a. Write an equation for the number of boxes you sell as a function of the price you charge. b. Denote "B" for number of boxes, and "P" for the price you charge. c. Assume the function is linear

OpenStudy (anonymous):

because it's linear: Change in boxes/change in price = slope 200-120/(3-4) = 80/-1=-80 B-200=-80(P-3)

OpenStudy (anonymous):

Thanks